40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 159     Entry time: Wed Jul 8 10:36:44 2009
Author: Michelle 
Type: Misc 
Category:  
Subject: Week 3 Update 

We have mirrors! And clamps! And posts! In short, everything we need to actually put all the optics in their proper places. Thus, this week I have successfully placed the steering mirrors and the high reflectance mirrors into their mounts without touching/breaking any of them, and soon I'll be clamping them in their proper locations.

Connor and I  found out that Aidan's laser isn't behaving itself. The manufacturer specifies the waist as being in a different location than where we calculated it to be. We sent the beam through a half-wave plate and then a polarizing beam splitter and then measured the power along each axis, and the ratios are strange. We can't get the beam to be entirely transmitted along one axis by rotating the half-wave plate, like we should be able to if the beam is coming out linearly polarized. So the hypothesis is...it's not coming out linearly polarized. It may be elliptical. For the time being we're going to split the beam from the other NPRO and use it for both experiments, at least while the alignment is taking place. In the case of the gyro, we'll actually be feeding back to the laser and altering its frequency, so we'll definitely need separate lasers for each experiment at that time.

The 495 mW NPRO's SOP has finally been (mostly) approved and it first lased on Monday. This laser's power supply is a bit weird - the display for the power it thinks the laser is putting out doesn't match at all with what the powermeter actually says the laser is putting out. So we're not paying attention to the power display. The drive current goes up to 1.00 A, and at this maximum current, the laser is putting out ~150 mW. Apart from not being able to get a higher power than this, the laser is fairly well behaved - the power fluctuations as measured by the powermeter are small. The results of output power vs. drive currentmeasurements are on the eLog.

Connor and I also measured the electronics noise and intensity noise at various different powers, and then did a theoretical shot noise calculation for each of these. The setup and procedure are described in more detail in a recent eLog entry, which also includes graphs and the matlab code used to generate them.

Over the next week I'll be setting up the optics table for a simple PDH lock using a fabry-perot cavity, aligning the laser, and setting up the proper feedback using the DAQ. (Which I'll need to talk to Dmass about).

Oh, and we found out yesterday that Connor and I can't do any hazardous tasks in the lab without supervision, which includes aligning the lasers, even at low power. Aidan and Alastair will both be gone from tomorrow until next week Friday, so this poses a slight challenge. Hopefully Dmass will be around?

That's about it for now, I think.

ELOG V3.1.3-